# DSwinIR: Rethinking Window-based Attention for Image Restoration

### v2

@@ IV-A Experimental Setup
TABLE III: Quantitative comparisons for Setting 2: five distinct degradation tasks. Results of our proposed DSwinIR are in bold. * denotes the results are adopted from previous work [15, 67].
@@
CBSD68 [79] SOTS [77] Rain100L [78] GoPro [80] LOL [81]
General SwinIR [9] ICCVW’21 30.59/0.868 21.50/0.891 30.78/0.923 24.52/0.773 17.81/0.723 25.04/0.835
MIRNet-v2 [91] TPAMI’22 30.97/0.881 24.03/0.927 33.89/0.954 26.30/0.799 21.52/0.815 27.34/0.875
Restormer [45] CVPR’22 31.49/0.884 24.09/0.927 34.81/0.962 27.22/0.829 20.41/0.806 27.60/0.881
NAFNet [34] ECCV’22 31.02/0.883 25.23/0.939 35.56/0.967 26.53/0.808 20.49/0.809 27.76/0.881
DRSformer* [13] CVPR’23 30.97/0.881 24.66/0.931 33.45/0.953 25.56/0.780 21.77/0.821 27.28/0.873
Retinexformer* [92] ICCV’23 30.84/0.880 24.81/0.933 32.68/0.940 25.09/0.779 22.76/0.834 27.24/0.873
FSNet* [35] TPAMI’23 31.33/0.883 25.53/0.943 36.07/0.968 28.32/0.869 22.29/0.829 28.71/0.898
MambaIR* [50] ECCV’24 31.41/0.884 25.81/0.944 36.55/0.971 28.61/0.875 22.49/0.832 28.97/0.901
All-in-One DL [88] TPAMI’19 23.09/0.745 20.54/0.826 21.96/0.762 19.86/0.672 19.83/0.712 21.05/0.743
TAPE [93] ECCV’22 30.18/0.855 22.16/0.861 29.67/0.904 24.47/0.763 18.97/0.621 25.09/0.801
Transweather [94] CVPR’22 29.00/0.841 21.32/0.885 29.43/0.905 25.12/0.757 21.21/0.792 25.22/0.836
AirNet [61] CVPR’22 30.91/0.882 21.04/0.884 32.98/0.951 24.35/0.781 18.18/0.735 25.49/0.846
IDR [62] CVPR’23 31.60/0.887 25.24/0.943 35.63/0.965 27.87/0.846 21.34/0.826 28.34/0.893
PromptIR* [15] NeurIPS’23 31.47/0.886 26.54/0.949 36.37/0.970 28.71/0.881 22.68/0.832 29.15/0.904
Gridformer* [89] IJCV’24 31.45/0.885 26.79/0.951 36.61/0.971 29.22/0.884 22.59/0.831 29.33/0.904
InstructIR [16] ECCV’24 31.40/0.887 27.10/0.956 36.84/0.973 29.40/0.886 23.00/0.836 29.55/0.907
AdaIR[63] ICLR’25 31.35/0.889 30.53/0.978 38.02/0.981 28.12/0.858 23.00/0.845 30.20/0.910
Perceive-IR [67] TIP’25 31.44/0.887 28.19/0.964 37.25/0.977 29.46/0.886 22.88/0.833 29.84/0.909
DSwinIR (Ours) 2025 31.34/0.885 30.09/0.975 37.77/0.982 29.17/0.879 22.64/0.843 30.19/0.913
@@ IV-A1 Evaluation Benchmarks and Metrics
Multi-Task All-in-One Restoration
This fundamental setting assesses the ability of a single model to serve as a versatile, general-purpose restorer. Following established conventions, we evaluate on two standard benchmarks. The three-task benchmark [61, 15] comprises Gaussian denoising, deraining, and dehazing. The more challenging five-task benchmark [63] extends this by incorporating motion deblurring and low-light enhancement, demanding greater generalization from a single set of model weights. The specific datasets for training and testing are detailed in Tab. I.
@@
This setting focuses on the practical and challenging domain of deweathering. We test DSwinIR’s robustness on both synthetic and real-world data. For the synthetic evaluation, we use the “AllWeather” collection [84], which includes images corrupted by snow, raindrops, and heavy rain, allowing for a controlled assessment across varied weather phenomena. To validate real-world efficacy, we employ a challenging benchmark [85] composed of datasets with authentic rain (SPA-Data [17]), snow (RealSnow [85]), and haze (REVIDE [86]), which feature complex, non-uniform artifacts that are the true test of a model’s generalization power.
@@
To push the boundaries of evaluation, we evaluate a composite degradation benchmark introduced in [95] to assess the model’s performance in more realistic and complex scenarios where multiple corruptions coexist. This benchmark is crucial for testing a model’s ability to disentangle and simultaneously mitigate interacting artifacts. We adopt a challenging test set by combining four fundamental degradation types: low-light (l), haze (h), rain (r), and snow (s). The evaluation protocol includes all single degradations, all six dual-degradation combinations (e.g., l+h, h+r), and two triple-degradation combinations (l+h+r, l+h+s). Success in this setting signals a higher level of architectural intelligence and robustness.
@@
While all-in-one models demonstrate versatility, a truly superior architecture must also excel when specialized. This experimental setting investigates the “specialization power” of DSwinIR by training it exclusively for a single task. We conduct focused evaluations on image deraining (Rain100L and the real-world SPA-Data) and dehazing (SOTS). Outperforming state-of-the-art specialist models in this context provides unequivocal evidence of our architecture’s foundational strength.
@@ IV-A2 Compared Methods
To comprehensively position DSwinIR within the literature, we conduct extensive comparisons against a spectrum of state-of-the-art models. This includes foundational, backbone-centric architectures such as Uformer [10], Restormer [45], DRSformer [13], and GridFormer [89], allowing for a direct benchmark of the architectural efficacy of our proposed DSwin attention. Furthermore, we contrast our model with cutting-edge prompt-based and instruction-guided all-in-one methods, including AirNet [61], PromptIR [15], InstructIR [16], and AdaIR [63]. Demonstrating superiority over this class of models highlights that our fundamental architectural innovations provide a more powerful and efficient solution than methods reliant on external semantic guidance or task-specific modules.
@@ IV-A3 Implementation Details
DSwinIR is trained end-to-end using the AdamW optimizer with an initial learning rate of $2\times 10^{-4}$, which is decayed using a cosine annealing schedule. Our model employs a four-level U-shaped architecture, with the number of DSwin Transformer blocks in the encoder stages set to [4, 6, 6, 8] from shallow to deep. We use a batch size of 8 and train on randomly cropped $\times 128$ 128 patches from the respective training images. All experiments are conducted using PyTorch on NVIDIA 3090 GPUs.
@@ IV-B Results and Analysis
In this section, we present a detailed analysis of DSwinIR’s performance across a hierarchy of benchmarks, starting from complex multi-degradation scenarios to specialized single-task settings. Our goal is to demonstrate not only the state-of-the-art capabilities of our model but also to dissect the sources of its power.
IV-B1 Dominance in All-in-One Image Restoration
@@
Following the classic benchmark setup [61, 15], we first evaluate on denoising, deraining, and dehazing. The quantitative results are presented in Tab. II. DSwinIR achieves a new state-of-the-art average PSNR of 32.72 dB, consistently outperforming all prior methods. This achievement is significant on two fronts. Firstly, it represents a substantial leap over other foundational architectures like GridFormer (32.19 dB) and DRSformer (31.26 dB), directly validating the efficacy of our DSwin attention in enhancing the intrinsic representation power of the network. Secondly, and perhaps more impressively, this performance is also superior to recent methods that employ additional, non-visual priors. For instance, DSwinIR surpasses the text-guided TextualDegRemoval [66] (32.63 dB) and the instruction-based InstructIR [16] (32.43 dB). That our vision-only backbone can outperform these multi-modal approaches strongly suggests that its architectural intelligence is sufficiently advanced to compensate for, and even exceed, the benefits of external guidance. The qualitative results in Fig. 4 corroborate these findings, showcasing DSwinIR’s superior ability to restore fine textures and remove complex artifacts.
@@
To further probe the model’s generalization capabilities, we turn to the more demanding five-task benchmark, which introduces deblurring and low-light enhancement. As detailed in Tab. III, DSwinIR again sets a new performance record with an average PSNR of 30.19 dB. The performance gains are particularly striking in tasks requiring significant spatial transformation, such as deraining (+0.93 dB over InstructIR) and dehazing (+2.99 dB over InstructIR). While its performance on low-light enhancement is competitive rather than leading, its exceptional strength across the other tasks secures the best overall average. This highlights the remarkable ability of the DSwinIR architecture to jointly learn from a highly diverse set of degradations within a single set of weights, without catastrophic interference and achieving a new state-of-the-art equilibrium.
TABLE IV: Quantitative comparisons for adverse weather removal including rain, haze, and snow. Missing values are denoted by ’–’.
Type Methods Venue Snow100K-S [82] Snow100K-L [82] Outdoor-Rain [84] RainDrop [83] Average
PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM
Task-Specific SPANet [17] CVPR’19 29.92 0.8260 23.70 0.7930
DesnowNet [82] TIP’18 32.33 0.9500 27.17 0.8983
HRGAN [96] CVPR’19 21.56 0.8550
MPRNet [6] CVPR’21 28.03 0.9192
AttentiveGAN [83] CVPR’18 31.59 0.9170
IDT [40] TIP’22 31.87 0.9313
NAFNet [34] ECCV’22 34.79 0.9497 30.06 0.9017 29.59 0.9027
Restormer [45] CVPR’22 36.01 0.9579 30.36 0.9068 30.03 0.9215 32.18 0.9408
All-in-One All-in-One [97] CVPR’20 28.33 0.8820 24.71 0.8980 31.12 0.9268 28.05 0.9023
TransWeather [94] CVPR’22 32.51 0.9341 29.31 0.8879 28.83 0.9000 30.17 0.9157 30.20 0.9094
Chen et al. [98] CVPR’22 34.42 0.9469 30.22 0.9071 29.27 0.9147 31.81 0.9309 31.43 0.9249
WGWSNet [99] CVPR’22 34.31 0.9460 30.16 0.9007 29.32 0.9207 32.38 0.9378 31.54 0.9263
WeatherDiff64 [100] TPAMI’23 35.83 0.9566 30.09 0.9041 29.64 0.9312 30.71 0.9312 31.57 0.9308
WeatherDiff128 [100] TPAMI‘23 35.02 0.9516 29.58 0.8941 29.72 0.9216 29.66 0.9225 31.00 0.9225
AWRCP [101] ICCV’23 36.92 0.9652 31.92 0.9341 31.39 0.9329 31.93 0.9314 33.04 0.9409
GridFormer [89] IJCV’24 37.46 0.9640 31.71 0.9231 31.87 0.9335 32.39 0.9362 33.36 0.9392
MPerceiver [102] CVPR’24 36.23 0.9571 31.02 0.9164 31.25 0.9246 33.21 0.9294 32.93 0.9319
DTPM [103] CVPR’24 37.01 0.9663 30.92 0.9174 30.99 0.9340 32.72 0.9440 32.91 0.9404
Histoformer [104] ECCV’24 37.41 0.9656 32.16 0.9261 32.08 0.9389 33.06 0.9441 33.68 0.9437
Figure 5: Vision results of AllWeather datasets, from up to bottom is samples for outdoor-rain, raindrop and snow test data. Our DSwinIR achieves better clarity even compared to the diffusion-based approaches.
IV-B2 Robustness in Deweathering Scenarios
Beyond general restoration, we investigate DSwinIR’s performance in the specific and practical domain of deweathering, which requires handling complex, real-world atmospheric phenomena.

### v3

@@ IV-A Experimental Setup
TABLE III: Quantitative comparisons for Setting 2: five distinct degradation tasks. * indicates numbers reported by prior work [PromptIR, zhang2024perceive].
@@
CBSD68 [martin2001database1] SOTS [li2019benchmarking] Rain100L [yang2017deep] GoPro [nah2017deep] LOL [wei2018deep]
General SwinIR [SwinIR] ICCVW’21 30.59/0.868 21.50/0.891 30.78/0.923 24.52/0.773 17.81/0.723 25.04/0.835
Uformer [Uformer] CVPR’22 30.49/0.883 23.40/0.919 33.32/0.962 26.26/0.807 21.16/0.817 26.93/0.877
MIRNet-v2 [MIRNet_v2] TPAMI’22 30.97/0.881 24.03/0.927 33.89/0.954 26.30/0.799 21.52/0.815 27.34/0.875
Restormer [Restormer] CVPR’22 31.49/0.884 24.09/0.927 34.81/0.962 27.22/0.829 20.41/0.806 27.60/0.881
NAFNet [chen2022simple] ECCV’22 31.02/0.883 25.23/0.939 35.56/0.967 26.53/0.808 20.49/0.809 27.76/0.881
DRSformer* [DRSformer] CVPR’23 30.97/0.881 24.66/0.931 33.45/0.953 25.56/0.780 21.77/0.821 27.28/0.873
Retinexformer* [Retinexformer] ICCV’23 30.84/0.880 24.81/0.933 32.68/0.940 25.09/0.779 22.76/0.834 27.24/0.873
FSNet* [CuiRCK24] TPAMI’23 31.33/0.883 25.53/0.943 36.07/0.968 28.32/0.869 22.29/0.829 28.71/0.898
MambaIR* [MambaIR] ECCV’24 31.41/0.884 25.81/0.944 36.55/0.971 28.61/0.875 22.49/0.832 28.97/0.901
VMambaIR [Shi2025VmambaIR] TCSVT’25 30.81/0.879 27.65/0.949 34.26/0.960 26.02/0.799 22.38/0.827 28.22/0.883
MaIR [Li2025MaIR] CVPR’25 30.98/0.882 27.72/0.950 34.04/0.946 25.68/0.791 22.17/0.821 28.11/0.876
All-in-One DL [dl] TPAMI’19 23.09/0.745 20.54/0.826 21.96/0.762 19.86/0.672 19.83/0.712 21.05/0.743
TAPE [TAPE] ECCV’22 30.18/0.855 22.16/0.861 29.67/0.904 24.47/0.763 18.97/0.621 25.09/0.801
Transweather [Transweather] CVPR’22 29.00/0.841 21.32/0.885 29.43/0.905 25.12/0.757 21.21/0.792 25.22/0.836
AirNet [AirNet] CVPR’22 30.91/0.882 21.04/0.884 32.98/0.951 24.35/0.781 18.18/0.735 25.49/0.846
IDR [IDR] CVPR’23 31.60/0.887 25.24/0.943 35.63/0.965 27.87/0.846 21.34/0.826 28.34/0.893
PromptIR* [PromptIR] NeurIPS’23 31.47/0.886 26.54/0.949 36.37/0.970 28.71/0.881 22.68/0.832 29.15/0.904
Gridformer* [Gridformer] IJCV’24 31.45/0.885 26.79/0.951 36.61/0.971 29.22/0.884 22.59/0.831 29.33/0.904
InstructIR [InstructIR] ECCV’24 31.40/0.887 27.10/0.956 36.84/0.973 29.40/0.886 23.00/0.836 29.55/0.907
AdaIR[cui2025adair] ICLR’25 31.35/0.889 30.53/0.978 38.02/0.981 28.12/0.858 23.00/0.845 30.20/0.910
Perceive-IR [zhang2024perceive] TIP’25 31.44/0.887 28.19/0.964 37.25/0.977 29.46/0.886 22.88/0.833 29.84/0.909
PoolNet [PoolNet] TIP’25 31.24/0.887 30.25/0.977 37.85/0.981 27.66/0.844 22.66/0.841 29.93/0.906
VLU-Net [Zeng2025VLUNet] CVPR’25 31.43/0.891 30.84/0.980 38.54/0.982 27.46/0.840 22.29/0.833 30.11/0.905
DSwinIR (Ours) 2025 31.34/0.885 30.09/0.975 37.77/0.982 29.17/0.879 22.64/0.843 30.20/0.913
@@ IV-A1 Evaluation Benchmarks and Metrics
Multi-Task All-in-One Image Restoration
This setting assesses the ability of a single model to serve as a versatile, general-purpose restorer. Following established conventions, we evaluate on two standard benchmarks. The three-task benchmark [AirNet, PromptIR] comprises Gaussian denoising, deraining, and dehazing. The more challenging five-task benchmark [cui2025adair] extends this by incorporating motion deblurring and low-light enhancement, demanding greater generalization from a single set of model weights. The specific datasets for training and testing are detailed in Table I.
@@
This setting focuses on the practical and challenging domain of deweathering. We test DSwinIR’s robustness on both synthetic and real-world data. For the synthetic evaluation, we use the “AllWeather” collection [test1_dataset_allweather], which includes images corrupted by snow, raindrops, and heavy rain, allowing for a controlled assessment across varied weather phenomena. To validate real-world efficacy, we employ a challenging benchmark [weather_data] composed of datasets with authentic rain (SPA-Data [SPANet_WangY0C0L19]), snow (RealSnow [weather_data]), and haze (REVIDE [Zhang2021Realhaze]), which feature complex, non-uniform artifacts that are the true test of a model’s generalization power. To further validate our method, we also conduct evaluations on the recent WeatherBench [Guan2025WeatherBench] dataset. This large-scale, real-world benchmark provides 41,402 training and 600 testing pairs. It presents a significant challenge by incorporating a diverse array of adverse weather conditions, including foggy, raining, and snowing scenarios captured during both daytime and nighttime.
@@
To push the boundaries of evaluation, we evaluate a composite degradation benchmark introduced in [guo2024onerestore] to assess the model’s performance in more realistic and complex scenarios where multiple corruptions coexist. This benchmark is crucial for testing a model’s ability to disentangle and simultaneously mitigate interacting artifacts. We adopt a challenging test set by combining four degradation types: low-light (l), haze (h), rain (r), and snow (s). The evaluation protocol includes all single degradations, all six dual-degradation combinations (e.g., l+h, h+r), and two triple-degradation combinations (l+h+r, l+h+s). Success in this setting signals a higher level of architectural intelligence and robustness.
@@
While all-in-one models demonstrate versatility, a truly superior architecture must also excel when specialized. This experimental setting investigates the “specialization power” of DSwinIR by training it exclusively for a single task. We conduct focused evaluations on image deraining (Rain100L and the real-world SPA-Data) and dehazing (SOTS). In addition, we employ further low-light evaluation benchmarks, including the real-world LOLv2 dataset. We also evaluate DSwinIR on a mixed task involving both low-light enhancement and deblurring on the LowBlur dataset, following the protocol from [Feijoo2025DarkIR]. Outperforming state-of-the-art specialist models in this context provides evidence of our architecture’s significant strength.
TABLE IV: Quantitative comparisons for adverse weather removal including rain, haze, and snow. Missing values are denoted by ‘—’.
Type Methods Venue Snow100K-S [Snow100K_dataset] Snow100K-L [Snow100K_dataset] Outdoor-Rain [test1_dataset_allweather] RainDrop [raindrop] Average
PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM
Task-Specific SPANet [SPANet_WangY0C0L19] CVPR’19 29.92 0.8260 23.70 0.7930
DesnowNet [Snow100K_dataset] TIP’18 32.33 0.9500 27.17 0.8983
HRGAN [outdoor] CVPR’19 21.56 0.8550
MPRNet [zamir2021multi] CVPR’21 28.03 0.9192
AttentiveGAN [raindrop] CVPR’18 31.59 0.9170
IDT [Xiao2023IDT] TIP’22 31.87 0.9313
NAFNet [chen2022simple] ECCV’22 34.79 0.9497 30.06 0.9017 29.59 0.9027
Restormer [Restormer] CVPR’22 36.01 0.9579 30.36 0.9068 30.03 0.9215 32.18 0.9408
All-in-One All-in-One [as2020] CVPR’20 28.33 0.8820 24.71 0.8980 31.12 0.9268 28.05 0.9023
TransWeather [Transweather] CVPR’22 32.51 0.9341 29.31 0.8879 28.83 0.9000 30.17 0.9157 30.20 0.9094
Chen et al. [ChenHTYDK22] CVPR’22 34.42 0.9469 30.22 0.9071 29.27 0.9147 31.81 0.9309 31.43 0.9249
WGWSNet [ZhuWFYGDQH23] CVPR’22 34.31 0.9460 30.16 0.9007 29.32 0.9207 32.38 0.9378 31.54 0.9263
WeatherDiff64 [weather_diffusion] TPAMI’23 35.83 0.9566 30.09 0.9041 29.64 0.9312 30.71 0.9312 31.57 0.9308
WeatherDiff128 [weather_diffusion] TPAMI‘23 35.02 0.9516 29.58 0.8941 29.72 0.9216 29.66 0.9225 31.00 0.9225
AWRCP [AWRCP_YeCBSXJYCL23] ICCV’23 36.92 0.9652 31.92 0.9341 31.39 0.9329 31.93 0.9314 33.04 0.9409
GridFormer [Gridformer] IJCV’24 37.46 0.9640 31.71 0.9231 31.87 0.9335 32.39 0.9362 33.36 0.9392
MPerceiver [AiHZW024] CVPR’24 36.23 0.9571 31.02 0.9164 31.25 0.9246 33.21 0.9294 32.93 0.9319
DTPM [DTPM_0001CCXQL024] CVPR’24 37.01 0.9663 30.92 0.9174 30.99 0.9340 32.72 0.9440 32.91 0.9404
Histoformer [Histoformer_SunRGWC24] ECCV’24 37.41 0.9656 32.16 0.9261 32.08 0.9389 33.06 0.9441 33.68 0.9437
Figure 6: Visual results on the AllWeather datasets; from top to bottom are samples for outdoor-rain, raindrop and snow test data. Our DSwinIR achieves better clarity even when compared to the diffusion-based approaches.
@@ IV-A2 Compared Methods
To comprehensively position DSwinIR within the literature, we conduct extensive comparisons against a spectrum of state-of-the-art models. This includes foundational, backbone-centric architectures such as Uformer [Uformer], Restormer [Restormer], DRSformer [DRSformer], GridFormer [Gridformer], MambaIR [MambaIR], VmambaIR [Shi2025VmambaIR], and MaIR [Li2025MaIR] allowing for a direct benchmark of the architectural efficacy of our proposed DSwin attention. Furthermore, we contrast our model with cutting-edge prompt-based and instruction-guided all-in-one methods, including AirNet [AirNet], PromptIR [PromptIR], InstructIR [InstructIR], AdaIR [cui2025adair], PoolNet [PoolNet], and VLUNet [Zeng2025VLUNet]. Demonstrating superiority over this class of models highlights that our DSwinIR provides a more powerful and efficient solution than methods reliant on external semantic guidance or task-specific modules.
@@ IV-A3 Implementation Details
DSwinIR is trained end-to-end using the AdamW optimizer with an initial learning rate of $2\times 10^{-4}$, which is decayed using a cosine annealing schedule. Our model employs a four-level U-shaped architecture, with the number of DSwin Transformer blocks in the encoder stages set to [4, 6, 6, 8] from shallow to deep. We use a batch size of 8 and train on randomly cropped $128\times 128$ patches from the respective training images. All experiments are conducted using PyTorch on 2 NVIDIA 3090 GPUs.
@@ IV-B Results and Analysis
In this section, we present a detailed analysis of DSwinIR’s performance across a hierarchy of benchmarks, ranging from complex multi-degradation scenarios to specialized single-task settings. Our aim is not only to establish the state-of-the-art performance of our model but also to provide insights into the architectural and mechanistic factors that contribute to its effectiveness. In the following result tables, Best and second-best results are highlighted.
IV-B1 Results on All‑in‑One Image Restoration
@@
Following the classic benchmark setup [AirNet, PromptIR], we first evaluate on denoising, deraining, and dehazing. The quantitative results are presented in Table II. DSwinIR achieves an average PSNR of 32.72 dB. This result demonstrates improvements over other foundational architectures like GridFormer (32.19 dB) and validates the efficacy of our DSwin attention. Furthermore, DSwinIR remains highly competitive even when compared to the latest prompt-based (e.g., AdaIR, 32.69 dB) and multimodal (e.g., VLUNet, 32.70 dB) methods. The ability of our vision-only backbone to perform at this level suggests its strong intrinsic representation power. The qualitative results in Figure 5 are consistent with these quantitative findings, showing that DSwinIR effectively restores fine textures and removes complex artifacts.
@@
To further probe the model’s generalization, we adopt the five-task benchmark, which introduces deblurring and low-light enhancement tasks. As detailed in Table III, DSwinIR achieves an average PSNR of 30.20 dB. This result is highly competitive with the top-performing methods, such as the prompt-based AdaIR (30.20 dB). The model’s performance is notably strong in tasks requiring significant spatial transformation, such as deraining (e.g., +0.93 dB over InstructIR) and dehazing (e.g., +2.99 dB over InstructIR). While its performance on low-light enhancement is competitive rather than leading, its overall strength across the other tasks secures this high average. This highlights the DSwinIR architecture’s ability to jointly learn from a diverse set of degradations within a single set of weights.
IV-B2 Results on Deweathering Datasets
Beyond general restoration, we investigate DSwinIR’s performance in the practical domain of deweathering, which requires handling complex, real-world atmospheric phenomena.